OpenStudy (anonymous):

minus 1 over 3 m-7=5

OpenStudy (anonymous):

-1/3m-7=5 -1/3m=12 m=-1/36

OpenStudy (anonymous):

I only have -36 on the multiple choice

OpenStudy (anonymous):

Can u draw the question

OpenStudy (anonymous):

same question as you wrote earlier

OpenStudy (anonymous):

here is another one 12+0.35x=20.05

OpenStudy (anonymous):

12+0.35x=20.05 0.35x=8.05 x=23

OpenStudy (anonymous):

should be m=-36